A Farmer has a 6% solution and a 12% solution of pesticide. How much of each must the farmer mix to have 300 gallons of an 8% solution for her spray?
Answer by TimothyLamb(4379)

pesticide solutions by volume:
s = (p / w)100
where
p = chemical volume
w = water volume
---
s = volume of strong solution (gal)
w = volume of weak solution (gal)
---
s + w = 300
s = 300 - w
12s + 6w = (8)300
---
12s + 6w = (8)300
12(300 - w) + 6w = (8)300
3600 - 12w + 6w = 2400
6w = 1200
---
w = 200 gal
s = 100 gal
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php